# The Vector Grounding Problem

### v2

@@ 7.2 The impact of fine-tuning
In a recent paper, Grindrod (2024) raises an objection to our argument from post-training. Grindrod acknowledges our claim that fine-tuning with human feedback introduces world-involving functions that could in principle establish the necessary historical relations for referential grounding. However, he objects:
The problem with this idea is that it would be odd to claim that a relatively marginal procedure that is designed to optimize the performance of a LLM on specific tasks ends up being the difference between intentionality and its lack. The reason, simply put, is that the fine-tuning procedures – whether RLHF, supervised training, or something else – are not the reason why we have such impressive LLMs today. (Grindrod 2024, p. 71)
@@
This objection misunderstands the purpose of our appeal to preference tuning. We are not claiming that preference tuning explains the impressive capabilities of LLMs or their ability to produce contextually appropriate outputs. Rather, we are specifically addressing how LLMs can overcome the Vector Grounding Problem – how their internal states can possess intrinsic meaning by being referentially grounded in the world. These are distinct questions.
Intrinsic meaning does not determine the complexity or impressiveness of outputs, neither in biological nor in artificial systems. Many relatively simple biological systems possess internal states with intrinsic meaning, while many complex systems might generate impressive outputs without intrinsic meaning. The quality of outputs and their status as intrinsically meaningful are orthogonal issues.
What matters for referential grounding is not how much a particular training procedure contributes to performance quality, but whether it establishes the right kind of functional relations between internal states and worldly entities. Even if preference tuning procedures make only marginal contributions to an LLM’s overall capabilities, they can fundamentally alter the normative conditions under which the model operates by introducing world-involving correctness standards.
@@ 7.3 Multimodality
The preceding discussion has identified two potential sources of referential grounding for LLMs: (1) preference tuning with human feedback that incorporates world-involving criteria such as factual accuracy, and (2) pre-training on next-token prediction, which may select for internal states with world-involving functional roles in limited domains where this is the most efficient way to achieve good prediction accuracy.
When we turn to multimodal and embodied systems, we need to examine whether and how these sources of referential grounding extend to such systems. Our analysis leads to a surprising conclusion: multimodality and embodiment are neither necessary nor sufficient conditions for referential grounding in artificial systems, though they can complement and potentially enhance referential grounding when the right kind of learning process is available.
Let us first consider a simple multimodal model like CLIP (Radford et al. 2021). CLIP is developed by jointly training an image encoder and a text encoder to predict the correct pairings of images and text captions using a contrastive learning objective. During training, CLIP receives batches of image-text pairs and learns to maximize the similarity between embeddings of matched pairs while minimizing the similarity between embeddings of unmatched pairs. This training approach allows CLIP to develop representations that capture semantic relationships between visual and linguistic content. Through contrastive pre-training with image-caption pairs, CLIP achieves what we called sensorimotor grounding, albeit to a very limited extent compared to biological organisms – namely connecting linguistic vector representations to pictorial vector representations. CLIP also arguably achieves some degree of relational grounding in the purely linguistic domain, although there is evidence that contrastive learning from image-caption pairs severely limits the degree to which this kind of model can induce syntactic structure compared to text-only LLMs (Yuksekgonul et al. 2023).
@@
As far as causal-informational relations go, the account we gave for LLMs should also apply to models like CLIP: since both images and their captions are shaped by human interactions with the world, CLIP should stand in complex, mediated causal-informational relations to the world. The images in CLIP’s training data were produced by humans using cameras to capture real-world scenes, and the captions were written by humans to describe those scenes. This establishes an indirect causal chain connecting CLIP’s internal representations to worldly entities and properties. However, the contrastive learning objective optimises purely for matching images with their captions, without establishing clear world-involving functional roles for CLIP’s internal states. There is no obvious selection pressure for internal states that track features of the world beyond what is necessary for the matching task, and no clear normative standard for success beyond the matching task itself. Thus, whether a model like CLIP can achieve referential grounding through the pathway outlined in the argument from pre-training is doubtful.
@@
By contrast with CLIP, some multimodal models are built with a pre-trained language model backbone. For example, typical VLMs that can receive images as well as text as input, such as Flamingo and LLaVA, leverage existing LLMs (with frozen weights) and learn a mapping between the visual encoder and the LLM (Bordes et al. 2024). This approach is computationally more efficient than training both vision and text encoders from scratch, as it only requires learning the connection between modalities. To the extent that the LLM backbone achieves referential grounding from pre-training on text alone, our framework entails that such model should inherit this property. The addition of a superficial mapping to the embedding space of a visual encoder provides a form of sensorimotor grounding, but it does not contribute to the referential grounding inherited from the pre-trained LLM.
@@ 7.4 Embodiment
Embodiment is often mentioned as the best way to secure that AI systems are grounded in the world, via their robotic bodies. Some embodied systems like SayCan (Ahn et al. 2022) integrate a pre-trained LLM with a robotic body in a way that keeps these components functionally separate. In SayCan, an LLM breaks down verbal instructions into sequences of ’atomic’ behaviours that the robot can enact through low-level visuomotor control. Importantly, the LLM component is not further trained or fine-tuned based on the success or failure of the robot’s actions. In such systems, the embodied aspect adds nothing to the referential grounding of the language model component. Without additional training that would establish new causal-historical relations between the LLM’s internal states and the outcomes of physical actions, the LLM’s capacity for referential grounding remains exactly what it was before integration with the robotic system. Like multimodality, physical embodiment alone does not automatically contribute to referential grounding. What matters is the causal-historical relationship established through training that selects for internal states with world-involving functional roles. Without such a learning history, embodiment remains external to the representational system itself.
When fine-tuning is involved, the story is a bit different. Recent advances in robotic learning have led to the development of sophisticated multimodal systems that integrate vision, language, and action. Models such as RT-2 (Zitkovich et al. 2023), RT-X (O’Neill et al. 2024), and OpenVLA (Kim et al. 2024) typically build upon language models or vision-language models that were pre-trained on text or multimodal data before being fine-tuned for robotic control. These models incorporate a pre-trained language model component and may inherit referential grounding from it. For instance, RT-2 builds upon vision-language models like PaLI-X and PaLM-E, which themselves incorporate language models pre-trained on text data. To the extent that these underlying language models achieve referential grounding through the mechanisms we have identified, the system as a whole may inherit this grounding.
However, VLA models go beyond merely inheriting referential grounding from their language model components. During embodied training, they are specifically trained to map visual inputs and natural language instructions to robotic actions that achieve specific goals in the physical world. This additional training process plausibly establishes new functional roles for the model’s internal states that are directly tied to successful physical interactions.
From our perspective, embodied training potentially enhances referential grounding by establishing or reinforcing causal-historical relations between internal states and features of the world that are relevant for successful manipulation. The training process selects for internal states that carry information about these features precisely because doing so improves performance on the physical tasks. This introduces an additional source of normativity tied to success in physical interaction, complementing whatever referential grounding was inherited from the pre-trained components.
@@ 8 Concluding Remarks
We have argued that the Symbol Grounding Problem, as introduced by Harnad (1990), can be adapted to apply to deep neural networks such as Generative Pre-Trained Models in we have called the Vector Grounding Problem. Despite initial intuitions, we have sought to demonstrate that at least a subset of current Generative Pre-Trained Models have the features necessary to be endowed with representations and outputs grounded in the world.
@@
While our discussion has focused on LLMs and, to a lesser extent, on their multimodal counterparts, any AI system fulfilling the causal-informational and historical relations requirements delineated above can possess representations with intrinsic meaning. Meeting these requirements is far from trivial, but nonetheless within reach of current AI models. As philosophical theories of representation have shown, in order to understand how cognitive systems – both biological and artificial – function, we must resist the temptation to assume that the presence of intrinsic meaning necessarily entails complex cognitive abilities such as understanding, knowledge, or consciousness. While having states imbued with intrinsic meaning might constitute one of the fundamental conditions for possessing these other cognitive abilities, it is insufficient on its own. Activations of specific populations of neurons in the mammalian cortex may represent the presence of edges in the visual scene, but these neuronal activations do not understand or possess knowledge about edges. We suggest that a similar claim can be made regarding the internal representations and outputs of LLMs and related connectionist models.
@@
We argued for this claim in two steps. First, we distinguished between various forms of grounding, which are often conflated in the cognitive science and computer science literature. We suggested that one particular form of grounding, referential grounding, is central to both the Vector Grounding Problem and the older Symbol Grounding Problem. Referential grounding pertains to the way in which representations hook onto the world. It is the most fundamental type of grounding, and arguably serves as the basis for all other notions of grounding (with the partial exception of relational grounding). Referential grounding requires two main factors: causal-informational relations between internal states and the world, and historical relations that imbue those states with functions to carry information about the world.
@@
Second, we have shown that certain LLMs trained exclusively on linguistic data, particularly those that undergo preference tuning, are referentially grounded. These models benefit from indirect causal-informational relations with worldly entities and properties, mediated by human interactions with the world, which in turn shape the linguistic inputs they receive during training and inference. Preference tuning introduces an additional learning objective that is directly world-involving, as it encompasses a requirement for truthfulness or factuality. The normativity of this requirement supplies an extra-linguistic accuracy standard for the representations and outputs of LLMs. More tentatively, we have suggested that even without preference tuning, the representations and outputs of LLMs may achieve referential grounding in some limited domains, through pre-training and in-context learning on specific kinds of input sequences, such as those describing game moves.
@@
This discussion led us to a counter-intuitive conclusion: some multimodal models trained on images in addition to text, like CLIP, are poor candidates for referential grounding, since they lack the kind of learning process that could supply world-involving functions. In fact, even some embodied systems that combine a ‘frozen’ LLM with a robotic body, like SayCan, also fail to acquire internal states and outputs that are referentially grounded. One benefit of our analysis is to highlight that multimodality and embodiment are neither necessary nor sufficient conditions for referential grounding in artificial systems, although they are certainly compatible with it if the right kind of learning process is available.

### v3

@@ 7.2 The impact of fine-tuning
In a recent paper, grindrodLargeLanguageModels2024 raises an objection to our argument from post-training. Grindrod acknowledges our claim that fine-tuning with human feedback introduces selection processes that could in principle establish the necessary historical relations for referential grounding. However, he objects:
The problem with this idea is that it would be odd to claim that a relatively marginal procedure that is designed to optimise the performance of a LLM on specific tasks ends up being the difference between intentionality and its lack. The reason, simply put, is that the fine-tuning procedures—whether RLHF, supervised training, or something else—are not the reason why we have such impressive LLMs today. (grindrodLargeLanguageModels2024, p. 71)
@@
This objection misunderstands the purpose of our appeal to preference tuning. We are not claiming that preference tuning explains the impressive capabilities of LLMs or their ability to produce contextually appropriate outputs. Rather, we are specifically addressing how LLMs can overcome the Vector Grounding Problem—how their internal states can possess intrinsic meaning by being referentially grounded in the world. These are distinct questions. Intrinsic meaning does not determine the complexity or impressiveness of outputs, neither in biological nor in artificial systems. Many relatively simple biological systems possess internal states with intrinsic meaning, while many complex systems might generate impressive outputs without intrinsic meaning. The quality of outputs and their status as intrinsically meaningful are orthogonal issues. What matters for referential grounding is not how much a particular training procedure contributes to performance quality, but whether it helps to establish the right kinds of relations between internal states and worldly entities. Even if preference tuning procedures make only marginal contributions to an LLM’s overall capabilities, they can fundamentally alter the normative conditions under which the model operates by introducing world-involving correctness standards.
@@ 7.3 Multimodality
At first blush, multimodal and embodied AI systems seem to be the most obvious candidates for overcoming the Vector Grounding Problem. Unlike pure LLMs, they receive sensory inputs (e.g., pixels) and/or interact physically with the world, so that the sort of ‘sensorimotor contact’ we discussed in subsection 4.2 seems to obtain. In section 5, we pointed out that an account of grounding based on mere causal contact is incomplete. Indeed, it follows from our analysis in section 6 that multimodality and embodiment are neither necessary nor sufficient for referential grounding, though they can be part of what makes referential grounding obtain when combined with the right kind of learning process.
@@
Let us first consider a simple multimodal model like CLIP (radfordLearningTransferableVisual2021). CLIP is developed by jointly training an image encoder and a text encoder to predict the correct pairings of images and text captions using a contrastive learning objective. During training, CLIP receives batches of image-text pairs and learns to maximise the similarity between embeddings of matched pairs while minimising the similarity between embeddings of unmatched pairs. As far as causal-informational relations go, the account we gave for LLMs should also apply to models like CLIP: since both images and their captions are shaped by human interactions with the world, CLIP should stand in complex, mediated causal-informational relations to the world. The images in CLIP’s training data were produced by humans using cameras to capture real-world scenes, and the captions were written by humans to describe those scenes. This establishes an indirect causal chain connecting CLIP’s internal states to worldly entities and properties.
However, the contrastive learning objective optimises purely for matching images with their captions, without establishing world-involving functions for CLIP’s internal states. There is no obvious selection pressure for internal states that track features of the world beyond what is necessary for the matching task, and no clear normative standard for success beyond the matching task itself. Consequently, CLIP lacks the selection history necessary to imbue its states with world-involving representational functions. It therefore fails to achieve referential grounding.
@@
By contrast with CLIP, some multimodal models are built with a pre-trained language model backbone. For example, typical VLMs that can receive images as well as text as input, such as Flamingo and LLaVA, leverage existing LLMs (with frozen weights) and learn a mapping between the visual encoder and the LLM (bordesIntroductionVisionLanguageModeling2024). This approach is computationally more efficient than training both vision and text encoders from scratch, as it only requires learning the connection between modalities. To the extent that the LLM backbone achieves referential grounding from pre-training on text alone, our framework entails that such a model should inherit this property. The addition of a superficial mapping to the embedding space of a visual encoder provides a form of sensorimotor grounding, but it does not contribute to the referential grounding inherited from the pre-trained LLM.
@@ 7.4 Embodiment
Embodiment is often mentioned as the best way to ensure that AI systems are grounded in the world, via their robotic bodies. Some embodied systems like SayCan (ahnCanNotSay2022) integrate a pre-trained LLM with a robotic body in a way that keeps these components functionally separate. In SayCan, an LLM breaks down verbal instructions into sequences of ‘atomic’ behaviours that the robot can enact through low-level visuomotor control. Importantly, the LLM component is not further trained or fine-tuned based on the success or failure of the robot’s actions—there is no selection of internal states on the basis of their contributions to successful behaviour. In such systems, the embodied aspect adds nothing to the referential grounding of the language model component. Without additional training that would establish new causal-historical relations between the LLM’s internal states and the outcomes of physical actions, the LLM’s capacity for referential grounding remains exactly what it was before integration with the robotic system. Like multimodality, physical embodiment alone does not automatically contribute to referential grounding.
When fine-tuning is involved, the story is a bit different. Recent advances in robotic learning have led to the development of sophisticated multimodal systems that integrate vision, language, and action. Models such as RT-2 (zitkovichRT2VisionLanguageActionModels2023a), RT-X (oneillOpenXEmbodimentRobotic2024), and OpenVLA (kimOpenVLAOpenSourceVisionLanguageAction2024) typically build upon language models or vision-language models that were pre-trained on text or multimodal data before being fine-tuned for robotic control. These models incorporate a pre-trained language model component and may inherit referential grounding from it. To the extent that the underlying language model achieves referential grounding through the mechanisms we have identified, the system as a whole may inherit this grounding.
However, VLA models go beyond merely inheriting referential grounding from their language model components. During embodied training, they are specifically trained to map visual inputs and natural language instructions to robotic actions that achieve specific goals in the physical world. This additional training process may endow the model’s internal states with world-involving, representational functions grounded in a history of learning through physical interactions with the world.
@@ 8 Concluding Remarks
We have argued that the Symbol Grounding Problem, as introduced by Harnad1990, also applies to deep neural networks such as generative pre-trained models, albeit in a slightly different form, which we have called the Vector Grounding Problem. We have sought to demonstrate that at least some current generative pre-trained models satisfy the requirements for having representations and outputs grounded in extra-linguistic referents.
@@
While our discussion has focused on LLMs and, to a lesser extent, on their multimodal counterparts, any AI system that fulfils the causal-informational and selection history requirements delineated above can possess representations with intrinsic meaning. Meeting these requirements is far from trivial, but nonetheless within reach of current AI models. As philosophical theories of representation have shown, to understand how cognitive systems—both biological and artificial—function, we must resist assuming that intrinsic meaning necessarily entails complex cognitive abilities such as understanding, knowledge, or consciousness. While having states imbued with intrinsic meaning might be a fundamental condition for possessing these cognitive abilities, it is insufficient on its own. Activations of specific populations of neurons in the mammalian cortex may represent the presence of edges in the visual scene, but these neuronal activations do not understand or possess knowledge about edges. We suggest that a similar claim can be made regarding the internal representations and outputs of LLMs and related connectionist models.
@@
We argued for this claim in two steps. First, we distinguished between various notions of grounding often conflated in the cognitive science and computer science literature. We suggested that one particular notion, referential grounding, is central to both the Vector Grounding Problem and the older Symbol Grounding Problem. Referential grounding pertains to how representations hook onto the world. It is the most fundamental type of grounding, and arguably serves as the basis for all other notions of grounding (with the partial exception of relational grounding). Referential grounding requires two main factors: causal-informational relations between internal states and the world, and a selection history that imbues those states with functions to carry information about the world.
@@
Second, we have shown that certain LLMs trained exclusively on linguistic data, particularly those that undergo preference tuning, are referentially grounded. These models benefit from indirect causal-informational relations with worldly entities and properties, mediated by human interactions with the world that shape the linguistic inputs they receive during training and inference. Preference tuning introduces an additional learning objective that is directly world-involving, as it encompasses a requirement for truthfulness or factuality. The normativity of this requirement supplies an extra-linguistic accuracy standard that selects and stabilises the internal states that contribute to meeting it. More tentatively, we suggested that even without preference tuning, the representations and outputs of LLMs may achieve referential grounding in limited domains through pre-training and in-context learning on specific input sequences, such as those describing game moves.
@@
This discussion led to a counter-intuitive conclusion: some multimodal models trained on images in addition to text, like CLIP, are poor candidates for referential grounding, since they lack the kind of learning process that could supply world-involving functions. Even embodied systems that combine a ‘frozen’ LLM with a robotic body, like SayCan, fail to acquire internal states and outputs that are referentially grounded. Our analysis highlights that multimodality and embodiment are neither necessary nor sufficient conditions for referential grounding in artificial systems, although they are certainly compatible with it given the right kind of learning process.